OpenStudy (anonymous):

Give the equivalent numerator if the denominator is (x - 2)(x - 1)(x - 6). 2/x-2

OpenStudy (anonymous):

The x-2 term is already there. So you need to multiply both the top and bottom by x-1 and by x-6 to get an equivalent fraction. \[\frac{2}{x-2}=\frac{2(x-1)(x-6)}{(x-2)(x-1)(x-6)}\]